# Tracing Galaxy Evolution in the Nearby Universe: The Role of Dark Matter

### v1

@@ III The Mass Estimations of LTGs
Given that LTGs have distinct structural components (stars/gas, bulge/disk), it is important to see how these different components affect the estimation of dynamical masses. In the case of the disk, we have assumed that the global behavior of the gas is governed by gravitational interaction. This assumption is not far from reality given that the force of gravity is the dominant force on large scales, keeping the gas within the galaxy and dictating its overall structure. The velocity dispersion used measures the gravitational effects on the gas. With this information, a mass estimation is made, which is calibrated using mass estimates obtained from rotation curves. We must remark that equation 4 provides an estimate of the dynamical mass under the assumption of virial equilibrium; however, this mass should not be confused with the virial mass associated with dark matter halos.
@@
Building on the previous discussion, we will offer definitions and discussion of these mass estimates, summarized in Table 1. In Table 1, the “Relevant Sections” column identifies the parts of the manuscript where the variables and biases linked to each dynamical mass estimate are described, particularly those we seek to mitigate.
@@
    where $a$ and $b$ represent the apparent semi-major and semi-minor axes of the galactic disk. The value 2.3548 corresponds to the conversion factor between the standard deviation and the full width at half maximum (FWHM) of a Gaussian profile, and is introduced to express the dispersion in terms of FWHM units. The inclination angle $i$ is estimated from the observed axial ratio ($b/a$), assuming an intrinsic flattening of 0.19 for edge-on late-type galaxies. This formulation follows the approach adopted by Mocz et al. (2012) where inclination corrections are essential for deriving intrinsic rotational velocities, surface brightness profiles, and dynamical parameters. The intrinsic axial ratio reflects the vertical thickness of galactic disks and is empirically calibrated from statistical samples of spiral galaxies. For more details on Equations 5 and 6, see Mocz et al. (2012).
@@
$\mathbf{M_{IA}}$ Variable (inclination angle) $H_{\alpha}$ gas A.1(b–c), A.7(c)
@@
Nigoche-Netro et al. (2022) developed a method to calibrate the masses of LTGs derived from Equation 4. The calibration was conducted using a subsample of LTGs that had rotational velocity data (Catinella et al., 2005). The calibration involved comparing the masses of the galaxies in the subsample obtained with Equation 4 (dynamical mass) with those obtained through rotational velocity measurements, using linear regressions. Their analysis considered various variables and sources of bias relevant to dynamical mass estimation, including magnitude, the Sérsic index $n$, concentration, colour, and galactic inclination angle. The key outcomes, as reported by Nigoche-Netro et al. (2022), are summarized below and presented in Table 2.
@@
  • When the magnitude is considered in the comparison of subsamples of LTGs, they find that the correlation coefficients attain the lowest values and the fit dispersions attain the highest values relative to the other variables considered in the analysis (Sérsic index $n$, concentration, colour, and angle of inclination). There are no significant differences found in the correlation fits or in the dispersion when comparing faint and bright galaxy samples (refer to their Figure 2 and Table 1). Therefore, the magnitude does not seem to influence the dynamical mass estimate. This paragraph addresses the variables and biases mentioned in Section A.8.
@@
  • When considering the Sérsic index $n$, the correlation is relatively high in all fits (see their Table 2). They observe that for $n\leq 0.8$, the slopes are lower than those for $n>0.8$. It is noticeable that for galaxies with $n\leq 0.8$, the correlation is lower when the dynamical mass is obtained using the stars compared to when it is obtained using the gas. For galaxies with $n>0.8$, the opposite occurs. It appears that morphology impacts the estimation of the dynamical mass. This paragraph considers the variables and biases discussed in Section A.1(b).
@@
  • When considering the LTGs concentration parameter (R90/R50), they find that for R90/R50 << 2.4, the slopes are lower than those for R90/R50 >> 2.4 (see their Table 3). They also find a relatively high correlation for all fits. In this case, they observe that the more concentrated LTGs (R90/R50 << 2.4) show a larger correlation when using the gas to estimate the dynamical mass, as opposed to using the stars. For the less concentrated galaxies (R90/R50 >> 2.4), the opposite occurs. The structure seems to affect the calculation of the dynamical mass. This paragraph considers the variables and biases discussed in Section A.1(b).
@@
  • When considering the colour of LTGs, they found that for blue galaxies ($g-r\leq 0.6$), the slopes are lower than those for red galaxies (gr>0.6g-r>0.6; see their Tables 4 and 5). They also find a relatively high correlation in all cases. It can be noticed that for blue galaxies, the correlation is lower when the dynamical mass is obtained using the stars compared to when it is obtained using the gas. For red galaxies, the opposite occurs. The determination of the dynamical mass seems to be influenced by the colour of the galaxy. This paragraph considers the variables and biases discussed in Section A.1(b).
@@
  • When taking into account the inclination angle of the galaxy, they observe that for smaller inclination angles, the slopes are greater than for larger inclination angles (see their Table 6). They also found a clear difference in correlation coefficient and dispersion for samples with smaller inclination angles compared to samples with larger inclination angles. The correlation of the fits is higher (indicating lower dispersion) for galaxies with inclination angles $i>66^{\circ}$, regardless of whether the velocity dispersion considered is that of the stars or the gas. The estimation of dynamical mass seems to be influenced by the galaxy’s inclination angle. This paragraph considers the variables and biases discussed in Section A.1(a).
@@
  • In general, the fits show a lower correlation (larger dispersion) when the galaxies are separated by luminosity values and a larger correlation (lower dispersion) when the galaxies are separated by inclination angle values (see their Tables 1-6 and Figures 2-8). In all cases, the best calibrations for the dynamic mass are those for which the inclination angle is relatively high ($i>66^{\circ}$). On the other hand, for LTG samples exhibiting traits such as relatively low Sérsic index, low concentration, or blue colour (indicating disk-dominated galaxies), the dynamical mass should be estimated using the gas velocity dispersion ($H_{\alpha}$ or the average velocity dispersion of the Balmer lines). In contrast, for cases with relatively high Sérsic index, high concentration, or red colour (suggesting bulge-dominated LTGs), it is more suitable to use the dynamical mass estimate derived from the velocity dispersion of the stars. The previous results are summarized in Table 2.
@@
Sérsic index ($n$) For $n\leq 0.8$ (disk-like), gas yields better fits; for Gas (low $n$), stars (high $n$)
@@
Concentration (R90/R50) R90/R50 << 2.4 (disk-like): better correlation with Gas (low conc.), stars (high conc.)
@@
Colour (grg-r) Blue galaxies (grg-r $\leq 0.6$, disk-like): gas preferred; Gas (blue), stars (red)
@@
Overall trend Inclination-based separation improves calibration; Gas (disk-like), stars (bulge-like)
tracer choice depends on morphology and colour.
@@ III.3 The stellar mass
Stellar population synthesis (SPS) methodologies are widely employed to determine the stellar masses of galaxies due to their more accurate and comprehensive estimations in comparison to alternative approaches. SPS relies on various components, including calculations of stellar evolution spanning from the main sequence to stellar death, spectral libraries of stars, empirical dust models, and stellar initial mass functions (Conroy, 2013, and references therein). These elements work in concert to transform the evolution of a diverse population of stars with varying metallicities and ages into a prediction for the evolving integrated light emitted by that stellar population. In the SDSS DR16 there are eight estimations of the total stellar mass obtained considering the elements mentioned before. In this study, we use these eight estimates, summarized in Table 3, and briefly describe them below. In Table 3, the “Relevant Sections” column identifies the parts of the manuscript where the variables and biases linked to each stellar mass estimate are described, particularly those we seek to mitigate.
@@
    $\mathbf{M_{ED}}$. StellarMassFSPSGranEarlyDust. Stellar masses for SDSS and BOSS galaxies were estimated using the Granada method, which involves fitting SPS models by Conroy et al. (2009) to SDSS photometry in the ugriz bands. The fitting process utilizes extinction-corrected model magnitudes scaled to the $i$-band c-model magnitude. This particular implementation of the method, referred to as the “early-star-formation” version, imposes a constraint on the epoch of star formation, restricting it to within 2 billion years after the Big Bang. In addition, it incorporates a dust extinction fitting procedure and adopts an initial mass function (IMF) of the form proposed by Van Dokkum (2008). This IMF is flexible with respect to a characteristic mass ($m_{c}$) and is continuous, thereby avoiding artificial discontinuities in the luminosity evolution of passive stellar populations. It closely resembles the widely used IMFs of Chabrier (2003) and Kroupa (2001) in the mass range around $1\mathbf{M_{Solar}}$, but allows $m_{c}$ to vary in order to explore possible temporal evolution of the IMF. Unlike the segmented power-law form of the Kroupa IMF, this continuous functional form provides a smoother and more realistic treatment of stellar population synthesis.
@@ IV Comparison of dynamical and stellar mass
As discussed earlier, the estimates of dynamical and stellar masses are subject to various biases. Specifically for dynamical mass, Nigoche-Netro et al. (2022) have observed that the most accurate calibrations for the dynamical masses of LTGs are linked to galaxies with relatively high inclination angles (see Section III.2). Their research highlights the importance of incorporating the morphology or colour of these galaxies through the use of specific variables, such as the Sérsic index ($n$), concentration index, or colour. Their findings indicate that for LTGs dominated by the disk, gas should be employed, while for LTGs dominated by the bulge, stars should be utilized. Therefore, in this study, we consider these findings to obtain the most accurate dynamical mass estimations for the galaxies of our galaxy sample. That is, according to the parameters of Nigoche-Netro et al. (2022), the dynamical masses calculated through the velocity dispersion of stars were obtained from subsamples of relatively high $n$, high concentration, and red LTGs with high inclination angles, while the dynamical masses calculated through the velocity dispersion of gas were obtained from subsamples of relatively low $n$, low concentration, and blue LTGs with high inclination angles.
@@
In this context, the term quasi-constant mass refers to logarithmic mass bins of width 0.1, while quasi-constant redshift corresponds to redshift bins of 0.01. This binning strategy significantly reduces the so-called geometric effect (Nigoche-Netro et al., 2015), which arises when broad distributions in mass and redshift produce artificial trends in the dynamical-to-stellar mass relation due to projection effects in the mass–redshift plane. By working in narrow intervals of both variables, one isolates the intrinsic behaviour of galaxies at fixed mass and redshift, allowing a cleaner investigation of how the difference between dynamical and stellar mass evolves with these parameters. Moreover, another strong reason to adopt narrow redshift ranges is that, within such intervals, sample completeness can be efficiently controlled, minimizing the impact of Malmquist bias (see section IV.1 for more details).
Applying this methodology, Nigoche-Netro et al. (2016) demonstrated that the discrepancy between dynamical and stellar masses for ETGs increases with dynamical mass but decreases with redshift. Such trends would be partially obscured or even misinterpreted if wide heterogeneous samples were used without binning. For this reason, in the present work we also adopt the approach of analyzing galaxy samples in quasi-constant mass and quasi-constant redshift intervals, ensuring that our conclusions about the relation between stellar, and dynamical mass are not dominated by geometric biases but instead reflect genuine astrophysical dependencies.
In the present analysis, we use the procedures of Nigoche-Netro et al. (2016) to analyze the relationship between dynamical and stellar mass as a function of dynamical mass and redshift for the LTG samples (see Sections A.10(a), and A.10(b)). We will compare the seven dynamical mass estimates with the eight stellar mass estimates, resulting in a total of 56 combinations. This examination will help us to determine whether the various variables and potential biases incorporated into these mass estimations have an impact on the outcomes related to dark matter.
@@
It is important to emphasize that, in order to avoid Malmquist bias—a selection effect whereby only the brightest objects are detected at larger distances—the comparison between samples must be restricted to the same mass intervals, and these intervals should correspond to ranges where the samples are complete, i.e., where all galaxies present in the Universe are represented (see the discussion of Figures 7 and 8 later in this same section). Completeness is evidenced by a well-populated distribution of galaxies across the considered mass and redshift ranges, a point that is crucial in the context of our analysis, since trends observed as a function of redshift or dynamical mass can otherwise be artificially driven by the sample’s incompleteness. By ensuring that the comparisons are made within complete mass intervals, we minimize selection effects and strengthen the reliability of the inferred correlations between dynamical and stellar masses. For a more detailed discussion of this point, see Section V.6.
@@
In Fig. 4, we present a mosaic showing the behavior of the stellar mass ($\mathbf{M_{ED}}$) as a function of redshift, assuming quasi-constant dynamical mass across the seven dynamical mass estimates. The logarithm of the dynamical mass ranges from approximately 10.3 (lower-left part of the panels) to 11.8 (upper-right part). For a fixed dynamical mass, the stellar mass systematically increases with redshift. Moreover, the slope of this increase depends on the mass regime: in the high-mass range, the stellar mass grows with redshift more gently than in the low-mass range. We also find that red, high Sérsic index $n$, and more concentrated LTGs—typically bulge-dominated—exhibit shallower slopes (see $\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, and $\mathbf{M_{IS}}$ in Figure 4), in contrast to blue, low $n$, and less concentrated LTGs—generally disk-dominated—which show steeper slopes (see $\mathbf{M_{KA}}$ and $\mathbf{M_{IA}}$ in Figure 4). The general observed pattern resembles that of ETGs reported by Nigoche-Netro et al. (2016) (see their Figure 6). In Appendix B, we display mosaics for the other stellar mass estimates ($\mathbf{M_{ND}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, $\mathbf{M_{PP}}$, $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, $\mathbf{M_{SF}}$). These confirm the same global trend: the difference between dynamical and stellar mass decreases with redshift. Therefore, the variables and possible biases discussed in Appendix A, which were considered in the different mass estimates, do not produce significant changes in the overall behavior. The only exceptions arise when dynamical masses are estimated using stars ($\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$) or gas ($\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$), indicating that structural components of galaxies are the main source of noticeable variations. Despite these differences, the overall trend remains robust: higher redshift corresponds to a smaller difference between dynamical and stellar mass.
@@
Figure 4: Behavior of stellar mass $({\bf M_{ED}})$ as function of redshift for quasi-constant dynamical mass. Each colour and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{ED}/{\bf M_{Solar}}})$ is approximately 0.062 dex.
@@
In Fig. 5, we show a mosaic of the difference between dynamical and stellar mass ($\mathbf{M_{ED}}$) as a function of dynamical mass, this time assuming quasi-constant redshift across the seven dynamical mass estimates. The trend lines at quasi-constant redshift are spaced in intervals of 0.01 in redshift. Redshift spans from approximately 0.025 (upper-left panels) to 0.225 (lower-right panels). The first column of the figure displays the behavior of the mass difference as a function of dynamical mass at quasi-constant redshift, covering the full redshift range available. To highlight the trends more clearly, the second column shows the same relation but restricted to three representative quasi-constant redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$). In this second column, we can see clearly that the trend lines are systematically offset, one below the other, indicating that at relatively low redshift the difference between dynamical and stellar mass is larger than at higher redshift (see also Figure 6). It also becomes evident that the evolution of the mass difference with dynamical mass at quasi-constant redshift is not the same for less massive galaxies (black dots $z\sim 0.025$) compared to more massive ones (dark green triangles $z\sim 0.165$). Specifically, the relation displays a saddle-like shape at low redshift, which changes slope when moving toward more massive galaxies and higher redshifts.

### v2

@@ III.1 The dynamical mass from the stars or gas of LTGs
Given that LTGs have distinct structural components (stars/gas, bulge/disk), it is important to see how these different components affect the estimation of dynamical masses. In the case of the disk, we have assumed that the global behavior of the gas is governed by gravitational interaction. This assumption is not far from reality, given that the force of gravity is the dominant force on large scales, keeping the gas within the galaxy and dictating its overall structure. The velocity dispersion used measures the gravitational effects on the gas. With this information, a mass estimation is made, which is calibrated using mass estimates obtained from rotation curves. We must remark that equation 4 provides an estimate of the dynamical mass under the assumption of virial equilibrium; however, this mass should not be confused with the virial mass associated with dark matter halos.
@@
Building on the previous discussion, we will offer definitions and a discussion of these mass estimates, summarized in Table 1. In Table 1, the “Relevant Sections” column identifies the parts of the manuscript where the variables and biases linked to each dynamical mass estimate are described, particularly those we seek to mitigate.
@@
    where $a$ and $b$ represent the apparent semi-major and semi-minor axes of the galactic disk. The value 2.3548 corresponds to the conversion factor between the standard deviation and the full width at half-maximum (FWHM) of a Gaussian profile, and is introduced to express the dispersion in terms of FWHM units. The inclination angle $i$ is estimated from the observed axial ratio ($b/a$), assuming an intrinsic flattening of 0.19 for edge-on late-type galaxies. This formulation follows the approach adopted by Mocz et al. (2012) where inclination corrections are essential for deriving intrinsic rotational velocities, surface brightness profiles, and dynamical parameters. The intrinsic axial ratio reflects the vertical thickness of galactic disks and is empirically calibrated from statistical samples of spiral galaxies. For more details on Equations 5 and 6, see Mocz et al. (2012).
@@
$\mathbf{M_{IA}}$ Variable (inclination angle) Hα gas A.1(b–c), A.7(c)
@@
Nigoche-Netro et al. (2022) developed a method to calibrate the masses of LTGs derived from Equation 4. The calibration was conducted using a subsample of LTGs that had rotational velocity data (Catinella et al., 2005). The calibration involved comparing the masses of the galaxies in the subsample obtained with Equation 4 (dynamical mass) with those obtained through rotational velocity measurements, using linear regressions. Their analysis considered various variables and sources of bias relevant to dynamical mass estimation, including magnitude, the Sérsic index $n$, concentration, color, and galactic inclination angle. The key outcomes, as reported by Nigoche-Netro et al. (2022), are summarized below and presented in Table 2.
@@
  • When the magnitude is considered in the comparison of subsamples of LTGs, they find that the correlation coefficients attain the lowest values and the fit dispersions attain the highest values relative to the other variables considered in the analysis (Sérsic index $n$, concentration, color, and angle of inclination). There are no significant differences found in the correlation fits or in the dispersion when comparing faint and bright galaxy samples (refer to their Figure 2 and Table 1). Therefore, the magnitude does not seem to influence the dynamical mass estimate. This paragraph addresses the variables and biases mentioned in Appendix A.8.
@@
  • When considering the Sérsic index $n$, the correlation is relatively high in all fits (see their Table 2). They observe that for $n\leq 0.8$, the slopes are lower than those for $n>0.8$. It is noticeable that for galaxies with $n\leq 0.8$, the correlation is lower when the dynamical mass is obtained using the stars compared to when it is obtained using the gas. For galaxies with $n>0.8$, the opposite occurs. It appears that morphology impacts the estimation of the dynamical mass. This paragraph considers the variables and biases discussed in Appendix A.1(b).
@@
  • When considering the LTGs concentration parameter (R90/R50), they find that for R90/R50 << 2.4, the slopes are lower than those for R90/R50 >> 2.4 (see their Table 3). They also find a relatively high correlation for all fits. In this case, they observe that the more concentrated LTGs (R90/R50 << 2.4) show a larger correlation when using the gas to estimate the dynamical mass, as opposed to using the stars. For the less concentrated galaxies (R90/R50 >> 2.4), the opposite occurs. The structure seems to affect the calculation of the dynamical mass. This paragraph considers the variables and biases discussed in Appendix A.1(b).
@@
  • When considering the color of LTGs, they found that for blue galaxies ($g-r\leq 0.6$), the slopes are lower than those for red galaxies (gr>0.6g-r>0.6; see their Tables 4 and 5). They also find a relatively high correlation in all cases. It can be noticed that for blue galaxies, the correlation is lower when the dynamical mass is obtained using the stars compared to when it is obtained using the gas. For red galaxies, the opposite occurs. The determination of the dynamical mass seems to be influenced by the color of the galaxy. This paragraph considers the variables and biases discussed in Appendix A.1(b).
@@
  • When taking into account the inclination angle of the galaxy, they observe that for smaller inclination angles, the slopes are greater than for larger inclination angles (see their Table 6). They also found a clear difference in correlation coefficient and dispersion for samples with smaller inclination angles compared to samples with larger inclination angles. The correlation of the fits is higher (indicating lower dispersion) for galaxies with inclination angles $i>66^{\circ}$, regardless of whether the velocity dispersion considered is that of the stars or the gas. The estimation of dynamical mass seems to be influenced by the galaxy’s inclination angle. This paragraph considers the variables and biases discussed in Appendix A.1(a).
@@
  • In general, the fits show a lower correlation (larger dispersion) when the galaxies are separated by luminosity values and a larger correlation (lower dispersion) when the galaxies are separated by inclination angle values (see their Tables 1-6 and Figures 2-8). In all cases, the best calibrations for the dynamic mass are those for which the inclination angle is relatively high ($i>66^{\circ}$). On the other hand, for LTG samples exhibiting traits such as relatively low Sérsic index, low concentration, or blue color (indicating disk-dominated galaxies), the dynamical mass should be estimated using the gas velocity dispersion ($H_{\alpha}$ or the average velocity dispersion of the Balmer lines). In contrast, for cases with relatively high Sérsic index, high concentration, or red color (suggesting bulge-dominated LTGs), it is more suitable to use the dynamical mass estimate derived from the velocity dispersion of the stars. The previous results are summarized in Table 2.
@@
Sérsic index ($n$) For $n\leq 0.8$ (disklike), gas yields better fits; for Gas (low $n$), stars (high $n$)
@@
Concentration (R90/R50) R90/R50 << 2.4 (disklike): better correlation with Gas (low conc.), stars (high conc.)
@@
color (grg-r) Blue galaxies (grg-r $\leq 0.6$, disklike): gas preferred; Gas (blue), stars (red)
@@
Overall trend Inclination-based separation improves calibration; Gas (disklike), stars (bulge-like)
tracer choice depends on morphology and color.
@@ III.3 The stellar mass
Stellar population synthesis (SPS) methodologies are widely employed to determine the stellar masses of galaxies due to their more accurate and comprehensive estimations in comparison to alternative approaches. SPS relies on various components, including calculations of stellar evolution spanning from the main sequence to stellar death, spectral libraries of stars, empirical dust models, and stellar IMFs (Conroy, 2013, and references therein). These elements work in concert to transform the evolution of a diverse population of stars with varying metallicities and ages into a prediction for the evolving integrated light emitted by that stellar population. In the SDSS DR16 there are eight estimations of the total stellar mass obtained considering the elements mentioned before. In this study, we use these eight estimates, summarized in Table 3, and briefly describe them below. In Table 3, the “Relevant Sections” column identifies the parts of the manuscript where the variables and biases linked to each stellar mass estimate are described, particularly those we seek to mitigate.
@@
    $\mathbf{M_{ED}}$. StellarMassFSPSGranEarlyDust. Stellar masses for SDSS and BOSS galaxies were estimated using the Granada method, which involves fitting SPS models by Conroy et al. (2009) to SDSS photometry in the ugriz bands. The fitting process utilizes extinction-corrected model magnitudes scaled to the $i$-band c-model magnitude. This particular implementation of the method, referred to as the “early-star-formation” version, imposes a constraint on the epoch of star formation, restricting it to within 2 billion years after the Big Bang. In addition, it incorporates a dust extinction fitting procedure and adopts an IMF of the form proposed by Van Dokkum (2008). This IMF is flexible with respect to a characteristic mass ($m_{c}$) and is continuous, thereby avoiding artificial discontinuities in the luminosity evolution of passive stellar populations. It closely resembles the widely used IMFs of Chabrier (2003) and Kroupa (2001) in the mass range around $1\mathbf{M_{Solar}}$, but allows $m_{c}$ to vary in order to explore possible temporal evolution of the IMF. Unlike the segmented power-law form of the Kroupa IMF, this continuous functional form provides a smoother and more realistic treatment of stellar population synthesis.
@@ IV Comparison of dynamical and stellar mass
As discussed earlier, the estimates of dynamical and stellar masses are subject to various biases. Specifically for dynamical mass, Nigoche-Netro et al. (2022) have observed that the most accurate calibrations for the dynamical masses of LTGs are linked to galaxies with relatively high inclination angles (see Section III.2). Their research highlights the importance of incorporating the morphology or color of these galaxies through the use of specific variables, such as the Sérsic index ($n$), concentration index, or color. Their findings indicate that for LTGs dominated by the disk, gas should be employed, while for LTGs dominated by the bulge, stars should be utilized. Therefore, in this study, we consider these findings to obtain the most accurate dynamical mass estimations for the galaxies of our galaxy sample. That is, according to the parameters of Nigoche-Netro et al. (2022), the dynamical masses calculated through the velocity dispersion of stars were obtained from subsamples of relatively high $n$, high concentration, and red LTGs with high inclination angles, while the dynamical masses calculated through the velocity dispersion of gas were obtained from subsamples of relatively low $n$, low concentration, and blue LTGs with high inclination angles.
@@
In this context, the term quasi-constant mass refers to logarithmic mass bins of width 0.1, while quasi-constant redshift corresponds to redshift bins of 0.01. This binning strategy significantly reduces the so-called geometric effect (Nigoche-Netro et al., 2015), which arises when broad distributions in mass and redshift produce artificial trends in the dynamical-to-stellar mass relation due to projection effects in the mass–redshift plane. By working in narrow intervals of both variables, one isolates the intrinsic behavior of galaxies at fixed mass and redshift, allowing a cleaner investigation of how the difference between dynamical and stellar mass evolves with these parameters. Moreover, another strong reason to adopt narrow redshift ranges is that, within such intervals, sample completeness can be efficiently controlled, minimizing the impact of Malmquist bias (see section IV.1 for more details).
Applying this methodology, Nigoche-Netro et al. (2016) demonstrated that the discrepancy between dynamical and stellar masses for ETGs increases with dynamical mass but decreases with redshift. Such trends would be partially obscured or even misinterpreted if wide heterogeneous samples were used without binning. For this reason, in the present work, we also adopt the approach of analyzing galaxy samples in quasi-constant mass and quasi-constant redshift intervals, ensuring that our conclusions about the relation between stellar and dynamical mass are not dominated by geometric biases but instead reflect genuine astrophysical dependencies.
In the present analysis, we use the procedures of Nigoche-Netro et al. (2016) to analyze the relationship between the dynamical and stellar mass as a function of dynamical mass and redshift for the LTG samples (see Appendices A.10(a), and A.10(b)). We will compare the seven dynamical mass estimates with the eight stellar mass estimates, resulting in a total of 56 combinations. This examination will help us to determine whether the various variables and potential biases incorporated into these mass estimations have an impact on the outcomes related to dark matter.
@@
It is important to emphasize that, in order to avoid Malmquist bias—–a selection effect whereby only the brightest objects are detected at larger distances—–the comparison between samples must be restricted to the same mass intervals, and these intervals should correspond to ranges where the samples are complete, i.e., where all galaxies present in the Universe are represented (see the discussion of Figures 7 and 8 later in this same section). Completeness is evidenced by a well-populated distribution of galaxies across the considered mass and redshift ranges, a point that is crucial in the context of our analysis, since trends observed as a function of redshift or dynamical mass can otherwise be artificially driven by the sample’s incompleteness. By ensuring that the comparisons are made within complete mass intervals, we minimize selection effects and strengthen the reliability of the inferred correlations between dynamical and stellar masses. For a more detailed discussion of this point, see Section V.6.
@@
In Fig. 4, we present a mosaic showing the behavior of the stellar mass ($\mathbf{M_{ED}}$) as a function of the redshift, assuming quasi-constant dynamical mass across the seven dynamical mass estimates. The logarithm of the dynamical mass ranges from approximately 10.3 (lower-left part of the panels) to 11.8 (upper-right part). For a fixed dynamical mass, the stellar mass systematically increases with redshift. Moreover, the slope of this increase depends on the mass regime; in the high-mass range, the stellar mass grows with redshift more gently than in the low-mass range. We also find that red, high Sérsic index $n$, and more concentrated LTGs—typically bulge-dominated—exhibit shallower slopes (see $\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, and $\mathbf{M_{IS}}$ in Figure 4), in contrast to blue, low $n$, and less concentrated LTGs—generally disk-dominated—which show steeper slopes (see $\mathbf{M_{KA}}$ and $\mathbf{M_{IA}}$ in Figure 4). The general observed pattern resembles that of ETGs reported by Nigoche-Netro et al. (2016, see their Figure 6). In Appendix B, we display mosaics for the other stellar mass estimates ($\mathbf{M_{ND}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, $\mathbf{M_{PP}}$, $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, $\mathbf{M_{SF}}$). These confirm the same global trend: the difference between dynamical and stellar mass decreases with redshift. Therefore, the variables and possible biases discussed in Appendix A, which were considered in the different mass estimates, do not produce significant changes in the overall behavior. The only exceptions arise when dynamical masses are estimated using stars ($\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$) or gas ($\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$), indicating that structural components of galaxies are the main source of noticeable variations. Despite these differences, the overall trend remains robust; a higher redshift corresponds to a smaller difference between dynamical and stellar mass.
@@
Figure 4: Behavior of stellar mass $({\bf M_{ED}})$ as a function of the redshift for quasi-constant dynamical mass. Each color and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30, while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{ED}/{\bf M_{Solar}}})$ is approximately 0.062 dex.
@@
In Fig. 5, we show a mosaic of the difference between dynamical and stellar mass ($\mathbf{M_{ED}}$) as a function of the dynamical mass, this time assuming quasi-constant redshift across the seven dynamical mass estimates. The trend lines at quasi-constant redshift are spaced in intervals of 0.01 in redshift. Redshift spans from approximately 0.025 (upper-left panels) to 0.225 (lower-right panels). The first column of the figure displays the behavior of the mass difference as a function of the dynamical mass at quasi-constant redshift, covering the full redshift range available. To highlight the trends more clearly, the second column shows the same relation but restricted to three representative quasi-constant redshifts: low (black dots, $z\sim 0.025$), intermediate (purple squares, $z\sim 0.095$), and high (dark green triangles, $z\sim 0.165$). In this second column, we can see clearly that the trend lines are systematically offset, one below the other, indicating that at relatively low redshift the difference between dynamical and stellar mass is larger than at higher redshift (see also Figure 6). It also becomes evident that the evolution of the mass difference with dynamical mass at quasi-constant redshift is not the same for less massive galaxies (black dots, $z\sim 0.025$) compared to more massive ones (dark green triangles, $z\sim 0.165$). Specifically, the relation displays a saddlelike shape at low redshift, which changes slope when moving toward more massive galaxies and higher redshifts.